### v1

# Assessing the Adversarial Security of Practical Perceptual Hashing Algorithms

Abstract
Perceptual hashing algorithms (PHAs) are utilized extensively for identifying illegal online content. Given their crucial role in sensitive applications, understanding their security strengths and weaknesses is critical. This paper compares three major PHAs deployed widely in practice: PhotoDNA, PDQ, and NeuralHash, and assesses their robustness against three typical attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks. Contrary to prevailing studies, this paper reveals that these PHAs exhibit resilience to black-box adversarial attacks when realistic constraints regarding the distortion and query budget are applied, attributed to the unique property of random hash variations. Moreover, this paper illustrates that original images can be reconstructed from the hash bits, raising significant privacy concerns. By comprehensively exposing their security vulnerabilities, this paper contributes to the ongoing efforts aimed at enhancing the security of PHAs for effective deployment.
Index Terms:
Perceptual hashing, security, privacy, blackbox adversarial attack, adversarial machine learning
I Introduction
Perceptual hashing, also known as robust hashing, is a content fingerprinting technique that generates similar binary hashes for perceptually similar contents and different hashes for different contents [1]. Perceptual hashing is useful in various applications such as content moderation, authentication, and detection. Numerous perceptual hash algorithms (PHAs) have been developed or deployed for practical applications [2]. An early example is the algorithm developed by Google and used on YouTube to identify copyrighted material [3]. Other examples include those used in online image search engines of Google, Microsoft Bing, etc [4].
PHAs have gained critical importance, particularly in responding to the reporting of illicit content. For instance, in cases involving Child Sexual Abuse Media (CSAM), US media service providers are legally obliged to examine suspected illegal content and report it to a “Cyber Tip Line” operated by the National Center for Missing and Exploited Children (NCMEC). Microsoft’s PhotoDNA and Meta’s PDQ have been developed for this purpose [5][6][7].
With a growing emphasis on user privacy, technologies like end-to-end encryption have become prevalent. This shift has led to the deployment of PHAs on the client side before data encryption [8]. For example, Apple has proposed to deploy NeuralHash on all its iOS devices [9]. However, deploying PHAs on the client side to scan client data has raised significant concerns regarding client data privacy, prompting widespread criticism about the robustness and privacy-preserving capabilities of the PHAs [10]. Especially, recent studies have almost unanimously criticized PHAs for their vulnerability to adversarial attacks and the potential leakage of information through hash bits [4][11][12].
Given their critical role in sensitive applications, it is imperative to thoroughly understand the security of PHAs. However, even for PHAs that have been widely deployed in the real world, their security remains insufficiently evaluated. Particularly concerning are new PHAs that are hastily deployed without adequate scrutiny. To address this gap, this paper systematically evaluates the adversarial robustness and privacy-preserving capabilities of three widely used PHAs: PhotoDNA, PDQ, and NeuralHash. To support this study, two new efficient attack algorithms are developed.
The major contributions of this paper are as follows.
  • A systematic comparison of the security and privacy-preserving capability of three major PHAs against three types of attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks, with the development of a query-efficient adversarial attack algorithm and a data-efficient hash inversion network.
  • Demonstrating the robustness of PHAs against malicious adversarial attacks when realistic constraints over distortion and query budget are imposed, contrary to previous studies that neglected such constraints and thus criticized PHAs as being not robust. The robustness is shown to be attributed to the unique property of random hash variations.
  • Revealing that the hash bits of all three PHAs leak significant information about the content of the original images, allowing for reliable reconstruction of the original image if sufficient context information is available.
The paper is organized as follows: Section II provides a literature review, Section III formulates the attacks and explains the robustness, Section IV presents the experiment results, and Section V concludes the paper. The source code of the experiments, including the binary executable of the PHAs, can be accessed at https://github.com/neddamj/nnhash.
II Literature Review
II-A Perceptual Hashing Algorithms
Among the vast amount of PHAs [2], this paper considers the following three major ones that have been deployed and used widely in practice:
PhotoDNA, released by Microsoft in 2009, has been used by over 70 companies including Cloudflare and Dropbox [6][13][14]. To create a hash, PhotoDNA breaks down an image into small pieces, analyzes each piece to extract its unique feature, and converts the features into hash bits. PhotoDNA generates a $144$-byte or $1152$-bit hash value for each input image.
PDQ, which stands for Perceptual algorithm utilising a Discrete Cosine Transform and outputting (amongst others) a Quality metric, was released by Meta in 2019 [7]. Inspired by pHash [15], PDQ uses the discrete cosine transform (DCT) to extract image features and outputs both a 256-bit hash value and a quality factor.
NeuralHash was released by Apple in 2021 with the intent to identify CSAM [9]. An image is first passed through a deep neural network (DNN) to produce a feature embedding. The embedding is then converted to a $96$-bit hash. The DNN is trained contrastively to ensure that the cosine similarity between perceptually similar images is high while the similarity of perceptually different images is low.
II-B Adversarial Machine Learning
The security of PHAs is commonly studied using the terminologies and algorithms of adversarial machine learning [16]. Specifically, Adversarial robustness measures the resilience of the machine learning models to adversarial attacks. Adversarial attacks can be classified as untargeted attacks and targeted attacks. In an untargeted attack, the objective is to slightly alter the input to prompt the model to produce a completely different output. Conversely, a targeted attack aims to slightly modify the input to elicit a specifically designed output.
According to the knowledge that the adversary has about the machine learning model, attacks can be categorized to whitebox attacks, where the adversary possesses complete knowledge of the model [16][17][18], and blackbox attacks, where the adversary only has access to the input and output of the model [19][20][21][22]. In black-box attacks, if the model’s outputs are continuous values like logits or probabilities, the attack is referred to as soft-label attack [19][20][21]. Otherwise, if the model’s outputs are solely hard-decision labels, the attack is known as a hard-label attack [22].
II-C Adversarial Robustness of PHAs
The adversarial robustness of conventional PHAs like pHash [2] was investigated in [23] under the whitebox setting, whereas the robustness under the blackbox setting was explored in [4]. Many of these studies suggested that PHAs were not robust to adversarial attacks.
The adversarial robustness of PhotoDNA, PDQ, and NeuralHash has not been adequately addressed, partly because they are new and partly because their source code is not publicly available. Regarding their robustness to normal image editing, [24] examined PhotoDNA, while [11] investigated PDQ and NeuralHash. Both studies demonstrated that PHAs were robust to certain edits like compression but vulnerable to others like cropping.
As far as we know, [25] was the first to study these PHAs’ robustness to adversarial attacks. It studied NeuralHash under the whitebox setting only, and showed that NeuralHash was not robust to whitebox attacks. [26] worked on PDQ and used the linearity of DCT to get adversarial images in a white-box setting. [27] generated adversarial images through linear interpolation between two given images.
Regarding blackbox attacks, [26] studied PDQ and claimed that it was not robust to untargeted blackbox attacks. [28] investigated PhotoDNA and PDQ, concluding that they were not robust to either targeted or untargeted attacks. These two studies closely resemble the current paper, yet this work differs from them in two key aspects. Firstly, previous works computed attack success rates without incorporating limitations on the adversary’s tolerance to image quality degradation or query budget. Such constraints are crucial in real-world scenarios. It is trivial that unconstrained blackbox attacks will always be successful, particularly in the untargeted setting. Secondly, none of the prior works conducted a comparative study of the three PHAs, nor did they investigate the challenge of hash inversion.
II-D Information Leakage from Hashes
Regarding privacy preservation, upon its release, Microsoft asserted that the hashes generated by PhotoDNA were irreversible and incapable of reconstructing the original images [29]. However, in a blog post [12], it was demonstrated that PhotoDNA hashes leak sufficient information about the original image to allow images to be synthesized from the hashes. To our knowledge, [12] is the sole study on reconstructing the original image from the hash, albeit focusing solely on PhotoDNA that has long hashes. It is unclear if PHAs with short hash suffer from this problem.
Another approach to investigate information leakage is to train a classifier to classify the image objects from the hash. [25][27] illustrated that NeuralHash hashes indeed revealed information about their source images by enabling the identification of the object’s class. Interestingly, these findings contradict those of [30] which examined PhotoDNA hashes and claimed that PhotoDNA was resistant to such classification attacks.
III Attack Algorithms
Consider a perceptual hash system where the service provider scans the client’s unencrypted content, extracts a hash for each image using a PHA, compares the hash with a database of known illicit hashes, and raises an alarm or forwards the unencrypted content for further analysis in case of hash match [8]. In this scenario, we consider two adversaries: the client adversary seeking to evade hash matching and the service provider (or some third party) adversary seeking to reconstruct the client’s original images from collected hashes.
We evaluate the security of PhotoDNA, PDQ, and NeuralHash against three types of attacks: benign image editing attacks and malicious adversarial blackbox attacks employed by the client adversary, and hash inversion attacks employed by the service provider adversary. Considering the client adversary, we assess the security of PHAs by measuring the adversary’s attack success rate, where a higher success rate indicates weaker security or robustness. Considering the service provider adversary, security is assessed through image reconstruction quality, where higher quality implies weaker security or privacy.
III-A Image Editing Attacks
Assume the original input to the PHA is an image ${\bf I}_{0}$, and the true output is an $N$-bit hash ${\bf h}_{0}={\cal H}({\bf I}_{0})$, where ${\cal H}(\cdot)$ denotes the hashing algorithm. The client adversary applies standard image editing operations over ${\bf I}_{0}$ to obtain an adversarial image ${\bf I}_{\rm adv}$ with a new $N$-bit hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$. A successful attack means ${\bf h}_{\rm adv}$ is sufficiently different from ${\bf h}_{0}$.
While prior works such as [11] and [24] have investigated image editing attacks on PHAs, a new evaluation is still necessary due to ongoing algorithm updates. Notably, NeuralHash has received at least one update since its first release. This study addresses the gap in knowledge by evaluating the security of the updated algorithms and, more importantly, by providing the first comparison of the three PHAs.
III-B Adversarial Blackbox attacks
III-B1 A Query-efficient Blackbox Attack Algorithm
This work focuses solely on blackbox, not whitebox, adversarial attacks, reflecting the practical scenario where PHA implementations are proprietary. For some PHAs, details regarding architecture and model weights are not publicly available due to their commercial nature. For instance, Microsoft enforces non-disclosure agreements for PhotoDNA users, and only an emulation is accessible. Similarly, while Ygvar et al. (2021) released weights for the first-generation NeuralHash, Apple has since updated the algorithm, keeping the new model confidential.
Existing blackbox attack algorithms, designed to operate on logits/probabilities or singular class decisions, cannot be directly applied to PHAs because PHAs output a limited number of hash bits, which are neither logits nor class labels. Instead, we propose a joint soft-label hard-label attack (JSHA) algorithm, where a soft-label attack step is applied first and its output is passed as input to a hard-label attack step. The objective of the soft-label attack step is to modify the image to change the hash without considering the image quality constraint, whereas the objective of the hard-label attack step is to reduce adversarial image quality degradation.
Consider the original image ${\bf I}_{0}$ with hash value ${\bf h}_{0}={\cal H}({\bf I}_{0})$. The objective of the client adversary is to generate an adversarial image ${\bf I}_{\rm adv}$ with new hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$. In an untargeted attack, ${\bf h}_{0}$ and ${\bf h}_{\rm adv}$ should be different enough. In a targeted attack, there is a target hash ${\bf h}_{\rm tar}\neq{\bf h}_{0}$, which may be obtained from some target image as ${\bf h}_{\rm tar}={\cal H}({\bf I}_{\rm tar})$, and we require that ${\bf h}_{\rm tar}$ and ${\bf h}_{\rm adv}$ are similar. In both attacks, the distance between ${\bf I}_{0}$ and ${\bf I}_{\rm adv}$ should be small.
To simplify notation, let us use normalized Hamming distance to evaluate the similarity of two $N$-bit hashes as
$D_{h}({\bf h}_{0},{\bf h}_{\rm adv})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}({\bf h}_{0}(n)={\bf h}_{\rm adv}(n))$ (1)
for two hashes with hash bits ${\bf h}_{0}(n)$ and ${\bf h}_{\rm adv}(n)$, where $\mathbb{I}(x)$ is the indicator function, i.e. $\mathbb{I}(x)=1$ if $x$ is true and 00 otherwise. In addition, we use per-pixel normalized $L_{2}$ norm to evaluate the distance of two images, i.e.
$L_{2}({\bf I}_{0},{\bf I}_{\rm adv})=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left({\bf I}_{0}(m)-{\bf I}_{\rm adv}(m)\right)^{2}}$ (2)
for two images with pixels ${\bf I}_{0}(m)$ and ${\bf I}_{\rm adv}(m)$. Other measures such as SSIM are also used in our experiments.
The first step of the proposed JSHA algorithm is to apply soft-label blackbox attack to generate an initial adversarial image ${\bf I}_{\rm init}$, with the objective to make the hash ${\bf h}_{\rm init}={\cal H}({\bf I}_{\rm init})$ far away from ${\bf h}_{0}$ for untargeted attacks or close to ${\bf h}_{\rm tar}$ for targeted attacks. This step is formulated as the optimization
$\max_{\delta}\;\;D_{h}\left({\bf h}_{0},{\cal H}({\bf I}_{0}+\delta)\right)$ (3)
in case of untargeted attacks, or
$\min_{\delta}\;\;D_{h}\left({\bf h}_{\rm tar},{\cal H}({\bf I}_{0}+\delta)\right)$ (4)
in case of targeted attacks. The optimization is conducted iteratively. During each iteration, the adversary queries the PHAs with its adversarial input ${\bf I}_{0}+\delta$ to get the hash, and uses the hash to optimize the loss (3) or (4). After reaching the query number threshold or the hash distance threshold, this first step stops with output ${\bf I}_{\rm init}={\bf I}_{0}+\delta$.
The second step applies a hard-label blackbox attack to reduce the image distortion while maintaining the hash distance obtained in the first step. Specifically, this step solves the following optimization
$\min\limits_{{\bf I}_{\rm adv}}\;\;L_{2}({\bf I}_{0},{\bf I}_{\rm adv}),\;\;\;\;\hskip 128.0374pt$ (5)
[…]
where $p$ is the pre-set hash distance threshold. Starting from ${\bf I}_{\rm adv}={\bf I}_{\rm init}$, the adversary queries the PHAs to get the hash and uses the queried hash to optimize (5) until the pre-set maximum query number or the threshold of $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ is reached.
The proposed algorithm is outlined in Algorithm 1. One of its major advantages is query efficiency. A large learning rate can be applied in the first step to get an initial adversarial sample quickly using a small number of queries. In the second step, since we start from ${\bf I}_{\rm init}$ which is only a noisy version of ${\bf I}_{0}$, the convergence is also fast. This makes our algorithm able to generate adversarial samples quickly with a very small number of queries.
Algorithm 1 JSHA: Joint Soft-Label and Hard-Label Attack
1:Initialization: Original input ${\bf I}_{0}$ and its hash ${\bf h}_{0}$; Target hash ${\bf h}_{\rm tar}$ in case of targeted attacks.
2:Step One: Unconstrained soft-label attack
  • The adversary uses queries to optimize (3) or (4)
  • Stop and output ${\bf I}_{\rm init}$ when the pre-set hash distance threshold or query budget is reached.
3:Step Two: Constrained hard-label attack
  • The adversary uses queries to optimize (5)
  • Stop and output ${\bf I}_{\rm adv}$ when the pre-set image distortion threshold or query budget is reached.
Refer to caption       Refer to caption
(a)                                (b)
Figure 1: (a) Hash inversion attack model. (b) Architecture of residual block.
III-B2 Robustness to Adversarial Blackbox Attacks
Algorithm 1 and most existing blackbox attack algorithms rely on gradient estimation to solve the optimization problem. Drawing on the property of PHAs and our prior work on adversarial robust analysis [31], we have the following proposition.
Proposition 1. Gradient estimation based on queried hash is not reliable.
Proof. Without loss of generality, consider the optimization problem (4). A common approach involves utilizing $K$ randomly perturbed images to query the PHAs and then using the queried hashes to estimate the gradient as follows
${\bf g}=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{\beta}{\bf u}_{k}D_{h}({\bf h}_{\rm tar},{\cal H}({\bf I}_{0}+\delta+\beta{\bf u}_{k}))$ (6)
where $\beta{\bf u}_{k}$ represents random perturbation [19]. Firstly, due to the nature of PHAs, overly small perturbations may not induce any change in hash. Hence, relatively large $\beta$ is required, but this renders the gradient estimation unreliable. Secondly, the hash ${\cal H}({\bf I}_{0}+\delta+\beta{\bf u}_{k})$ randomly varies with the small perturbation $\beta{\bf u}_{k}$. The random variation can be modeled as a noise term $Z$ with variance $\sigma^{2}$. Following the analysis of [31], this noise is amplified by $1/\beta$, leading to the gradient estimate ${\bf g}$ being affected by a noise power $\sigma^{2}/\beta^{2}$, hence making it unreliable.
Due to the random nature of PHAs, it is almost impossible to limit the change of hash to just one or a few hash bits by perturbations. Practical $\sigma$ is usually around $0.1$. $\beta$ may be around $0.001$. As per [31], the signal-to-noise-ratio (SNR) of the gradient estimation can fall well below $-40$ dB, which makes the adversary’s optimization not convergent or converge extremely slowly. This implies that PHAs are theoretically robust to blackbox attacks under the query budget constraint.
III-C Hash Inversion Attacks
GAN (generative adversarial network) was used by [12]. However, the service provider adversary may need more efficient hash inversion models because the training data may not be abundant in practice. For this, we train new and more efficient generative models to reconstruct the original input images from the hashes. We designed a decider style DNN, similar to the Fast Style Transfer [32], to convert the $N$-bit hashes to $(64,64)$ images. This DNN is mainly a combination of residual blocks, which are described in Fig. 1(b), with fractionally-strided convolutions/deconvolutions that are used for learned upsampling. The entire architecture is described in Fig. 1(a).
For a given training dataset, we first compute the hashes of each image offline and store the image-hash pairs. We then train the model outlined above to reconstruct the images as closely as possible by minimizing the mean-square-error (MSE) loss between the DNN output and the image from which the hash was generated.
IV Experiment
IV-A Performance Metrics
For assessing the robustness against blackbox attacks, we employ the attack success rate as the evaluation metric. Given a set of $M$ original images ${\bf I}_{0}^{m}$ with their corresponding hashes ${\bf h}_{0}^{m}$, the adversary generates adversarial images ${\bf I}_{\rm adv}^{m}$ along with their hashes ${\bf h}_{\rm adv}^{m}$. The attack success rate is defined as
${\rm ASR}(p)=$
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{0}^{m},{\bf h}_{\rm adv}^{m})>p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (7)
for untargeted attacks, and
${\rm ASR}(p)=$
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{\rm tar}^{m},{\bf h}_{\rm adv}^{m})<p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (8)
for targeted attacks (with the target hash ${\bf h}_{\rm tar}$). $p\in[0,1]$ is the pre-set hash distance threshold and $\theta$ is the pre-set image distortion threshold. The adversary can adjust the threshold $p$ to control the tradeoff between the true positive rate (TPR) and the false positive rate (FPR) of the hash or image similarity decisions.
Untargeted attacks use a large $p$ because ${\bf h}_{\rm adv}$ should be different from ${\bf h}_{0}$. As shown in Fig. 2, when we calculated $D_{h}({\bf h}_{0},{\bf h}_{1})$ of distinct images of the ImageNet validation dataset, we found that very few pairs of images had $D_{h}({\bf h}_{0},{\bf h}_{1})<0.3$, which was also observed in [11]. Therefore, we primarily use $p=0.3$ in (7) to determine the success of untargeted attacks. In contrast, targeted attacks use a small $p$ to make sure ${\bf h}_{\rm adv}$ is sufficiently close to ${\bf h}_{\rm tar}$. Prior studies [11] indicate that many normal image edits would change an image’s hashes to the level of $D_{h}({\bf h}_{0},{\bf h}_{1})<0.1$. Therefore, we mainly use $p=0.1$ in (8) for assessing the success of targeted attacks.
To comprehensively assess attack success, we collect ASR$(p)$ for all $p=0.1$, $0.2$, $0.3$, $0.4$ during experiments. We set the image distortion threshold $\theta=0.1$.
Figure 2: Cumulative distribution function (cdf) of normalized hamming distance $D_{h}({\bf h}_{0},{\bf h}_{1})$ between hashes of distinct images in the ImageNet validation dataset.
For hash inversion attacks, we use the $L_{2}$ distortion (2), the SSIM (Structural Similarity Index Measure) [33], and the LPIPS (Learned Perceptual Image Patch Similarity) [34] between the reconstructed image and the true image to evaluate the performance. No ASR or query budget is involved.
IV-B Experiment Results of Image Editing Attacks
We experimented with four image editing operations that can keep the original high image quality: JPEG compression with random quality factors, resizing images with random scaling factors, filtering images with the vignette filter, or rotating images to random degrees. We calculated the hashes ${\bf h}_{0}$ and ${\bf h}_{1}$ of the images before and after the editing, and derived their distance $D_{h}({\bf h}_{0},{\bf h}_{1})$. Using 1000 images from the ImageNet validation dataset, we obtained ASR$(p)$ for various $p$ based on (7), wherein the $L_{2}$ condition was removed by setting $\theta$ to infinity.
Experiment results are presented in Tables I. The data for ASR(.3) is highlighted in black because $p=0.3$ is the threshold we suggested to focus on. It is evident that all three PHAs were robust to compression and resizing, but not to filtering and rotation. In conjunction with the findings in [11][25], the PHAs are robust to many standard image editing operations except the following:
  • PhotoDNA is not robust to: filtering, rotating, resizing, mirroring, bordering, cropping;
  • PDQ is not robust to: filtering, rotating, mirroring [11], bordering [11], cropping [11];
  • NeuralHash is not robust to: filtering, rotating, mirroring [11].
NeuralHash was the most robust among the three. Surprisingly, the newest edition of NeuralHash that we experimented with was still not robust to filtering, rotating, and mirroring, similar to the first edition experimented in [11][25], even though DNNs can be trained to be invariant to such standard operations with data augmentation.
TABLE I: Adversary’s ASR($p$) with Image Editing Attacks
ASR Compress Resize Filter Rotate
PhotoDNA ASR(.1) 3% 99% $100\%$ $100\%$
PhotoDNA ASR(.2) $0\%$ 55% $100\%$ $100\%$
PhotoDNA ASR(.3) 0% 0% 94% 100%
PhotoDNA ASR(.4) $0\%$ 0% $37\%$ $98\%$
PDQ ASR(.1) 0% 27% 88% 100%
PDQ ASR(.2) 0% 0% 67% 100%
PDQ ASR(.3) 0% 0% 44% 100%
PDQ ASR(.4) 0% 0% 27% 100%
NeuralHash ASR(.1) 0% 22% 96% 100%
NeuralHash ASR(.2) 0% 0% 82% 100%
NeuralHash ASR(.3) 0% 0% 57% 96%
NeuralHash ASR(.4) 0% 0% 12% 31%
IV-C Experiment Results of Untargeted Blackbox Attacks
In implementing our proposed JSHA algorithm, we modified three different soft-label attack algorithms: SimBA [21], NES [19], and ZOsignSGD [20], to use in Step One, and modified the hard-label attack algorithm HSJA [22] to use in Step Two. For instance, “SimBA+HSJA” means our JSHA with a modified SimBA as Step One and a modified HSJA as Step Two. Modifications included removing/changing constraints and adjusting hyper-parameters, etc. Specifically, in Step One, we skipped image distortion constraints and adopted extremely large perturbation step sizes, e.g, using a perturbation magnitude of $\epsilon=0.9$ for SimBA instead of the original $\epsilon=0.01$ as described in [21].
The experiments were conducted on 100 randomly selected images from the ImageNet validation dataset. Step One was executed for a total of 3000 queries unless the early stopping criterion was met. Step Two was run for an additional 2000 queries unless meeting the early stopping criterion. The main challenge for us was the CPU-based non-parallel PHA executables, which limited the number of images or queries we could process. In comparison, [28] only utilized 30 images in experiments.
Experiment results are presented in Tables II. We compared the proposed JSHA with the straightforward application of standard SimBA [21], NES [19], and ZOsignSGG [20] algorithms. We also compared it with [26] and [28].
With a focus on ASR(0.3), the data indicated that NeuralHash was relatively robust because the maximum ASR$(0.3)$ was $28\%$ only. PDQ was robust against all adversarial attacks, with a maximum ASR$(0.3)$ of 1% only. It’s worth noting that although [26] achieved a much higher attack success rate against PDQ, it required significantly more queries. It could afford millions of queries due to its access to the PDQ source code, enabling faster computation.
Furthermore, the experiment data revealed that our proposed JSHA outperformed standard blackbox attack algorithms. SimBA+HSJA surpassed [28] in attacking PhotoDNA, as the latter produced more noisy adversarial images and utilized more queries.
Sample adversarial images generated with our proposed JSHA with various pre-set threshold $p$ are depicted in Fig. 3, with their corresponding distortion metrics listed in Table III. Notably, the three images with $p=0.4$ were deemed invalid adversarial images as their distortion was not under the threshold $\theta=0.1$, resulting in extremely noisy images. Similarly, for PDQ, the adversarial images with $p=0.2$ and $0.3$ were also invalid.
In summary, contrary to previous studies suggesting that PHAs were not robust against untargeted attacks, we found that some PHAs like PDQ and NeuralHash could be considered as robust if realistic constraints regarding image distortion and query budget are taken into account.
TABLE II: Adversary’s ASR($p$) with Untargeted Black-box Attacks.
(PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
ZO- ZOsign
SimBA NES+ sign- -SGD
ASR SimBA +HSJA NES HSJA SGD +HSJA
PH ASR(.1) $100\%$ $100\%$ $18\%$ $100\%$ $1\%$ $98\%$
PH ASR(.2) $80\%$ $100\%$ $2\%$ $98\%$ $0\%$ $90\%$
PH ASR(.3) $1\%$ 92% $0\%$ $69\%$ $0\%$ $19\%$
PH ASR(.4) $0\%$ $0\%$ $0\%$ $1\%$ $0\%$ $0\%$
PD ASR(.1) 0% 45% 0% 64% 0% 8%
PD ASR(.2) 0% 1% 0% 10% 0% 0%
PD ASR(.3) 0% 0% 0% 1% 0% 0%
PD ASR(.4) 0% 0% 0% 0% 0% 0%
NH ASR(.1) 4% 88% 0% 99% 0% 94%
NH ASR(.2) 2% 23% 0% 77% 0% 42%
NH ASR(.3) 0% 14% 0% 28% 0% 2%
NH ASR(.4) 0% 0% 0% 0% 0% 0%
[28] PH ASR(.3)$=$ 90%, ($L_{2}=0.2$, HSJA with $3\times 10^{4}$ queries)
[26] PD ASR(.3)$=$ 73%, ($L_{2}=0.1$, up to $8\times 10^{6}$ queries)
(a) Original image ${\bf I}_{0}$.
PhotoDNA                  PDQ                   NeuralHash
(b) Adversarial images ${\bf I}_{\rm adv}$.
Figure 3: Samples of an original image and the adversarial images created by the proposed untargeted blackbox attack algorithm JSHA (NES+HSJA).
TABLE III: Distortion of the Adversarial Image Samples in Fig. 3(b).
Measure Threshold $p$ PhotoDNA PDQ NeuralHash
$p=0.1$ 0.02 0.07 0.02
$L_{2}$ $p=0.2$ 0.02 0.11 0.04
$p=0.3$ 0.05 0.14 0.07
$p=0.4$ 0.17 0.19 0.19
IV-D Experiment Results of Targeted Blackbox Attacks
The experiment setting of targeted attacks was similar to that of untargeted attacks, except for utilizing ${\bf h}_{\rm tar}$ obtained from random images instead of ${\bf h}_{0}$. Results are presented in Tables IV. With a focus on ASR(0.1) in this case, we observed that all targeted attacks failed, yielding ASR$(0.1)=0$. To our knowledge, no literature has ever reported non-zero ASR$(0.1)$, suggesting that all the PHAs are robust against targeted attacks.
The robustness is extremely strong because of ASR$(p)\approx 0$ across all $p$. The only exception was NeuralHash, which exhibited ASR(0.4) of 43% with SimBA. However, this isn’t indicative of successful adversarial attacks; rather, it’s due to numerous hashes generated by NeuralHash having $D_{h}({\bf h}_{0},{\bf h}_{1})$ as low as 0.3, as illustrated in Fig. 2. Under NeuralHash, many images are inherently adversarial when evaluated at $p=0.4$.
In summary, contrary to the prior claims that PHAs were not robust against targeted attacks [28], our findings demonstrate the strong robustness of all PHAs against targeted blackbox attacks. This is significant to the practical application of PHAs because the client adversary needs targeted attacks to ensure that its adversarial hashes do not inadvertently match illicit hashes in the database. The strong robustness means it is hard for the adversary to undertake either untargeted or targeted attacks without being caught.
TABLE IV: Adversary’s ASR($p$) with Targeted Blackbox Attacks.
(PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
ZO- ZOsign
SimBA NES+ sign- -SGD
ASR SimBA +HSJA NES HSJA SGD +HSJA
PH ASR(.1) 0% 0% 0% 0% 0% 0%
PH ASR(.2) 0% 0% 0% 0% 0% 0%
PH ASR(.3) 0% 0% 0% 0% 0% 0%
PH ASR(.4) 1% 2% 0% 0% 0% 0%
PD ASR(.1) 0% 0% 0% 0% 0% 0%
PD ASR(.2) 0% 0% 0% 0% 0% 0%
PD ASR(.3) 0% 0% 0% 0% 0% 0%
PD ASR(.4) 0% 0% 0% 0% 0% 0%
NH ASR(.1) 0% 0% 0% 0% 0% 0%
NH ASR(.2) 0% 0% 0% 0% 0% 0%
NH ASR(.3) 0% 0% 0% 0% 0% 0%
NH ASR(.4) 43% 11% 0% 0% 0% 0%
[28] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
[28] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
IV-E Experiment Results of Hash Inversion Attacks
To demonstrate the service provider adversary’s capability to reconstruct original images from their hashes under some contextual knowledge, such as the group of images to which the original images may belong, we conducted experiments using two datasets MNIST and CelebA as known contexts. For each hashing algorithm and dataset, we trained the inversion network depicted in Fig. 1 for 50 epochs, with a batch size of 64 and an initial learning rate of 0.005. We utilized the MSE loss function with the AdamW optimizer and the cosine annealing learning rate scheduler.
Results are presented in Table V. Samples of synthesized images along with their original images can be observed in Fig. 4 and Fig. 5. It is evident that the hash leaks sufficient information for image reconstruction when certain contextual information is known, such as whether the image belongs to MNIST digits or CelebA photos. Although no image quality metrics were given in [12], Fig. 5(c) shows that our proposed model had a performance comparable to that of [12].
In Table V, smaller $L_{2}$ and LPIPS mean better image reconstruction, while larger SSIM means better image reconstruction. PhotoDNA has the longest hashes at 1152-bits, PDQ has 256-bit hashes and NeuralHash has 96-bit hashes. Intuitively, this implies that PhotoDNA hashes contain the most information about the original images while NeuralHash hashes contain the least. Table V supported this intuition.
The results in the MNIST for PDQ were left blank because PDQ could not work on MNIST. PDQ outputs quality scores of 0 for every image, and each of the hash values calculated turned out to be a 256-dimensional vector of zeros for MNIST images.
TABLE V: Quality of Adversarial Images Created in Hash Inversion Attacks (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
Dataset Hash $L_{2}$ SSIM LPIPS
PH $\mathbf{0.08}\pm 0.02$ $\mathbf{0.79}\pm 0.02$ $\mathbf{0.18}\pm 0.03$
MNIST PD - - -
NH $0.19\pm 0.05$ $0.35\pm 0.10$ $0.33\pm 0.08$
PH $\mathbf{0.13}\pm 0.03$ $\mathbf{0.57}\pm 0.08$ $\mathbf{0.40}\pm 0.07$
CelebA PD $0.23\pm 0.07$ $0.44\pm 0.07$ $0.43\pm 0.05$
NH $0.23\pm 0.05$ $0.29\pm 0.09$ $0.53\pm 0.07$
(a) PhotoDNA. Left: (.14, .83, .05). Right: (.12, .86, .07)
(b) NeuralHash. Left: (.26, .61, .10). Right: (.34, .52, .14)
Figure 4: Samples of true images and adversarial images generated by hash inversion attacks based on hashes of (a) PhotoDNA, and (b) NeuralHash. The numbers are ($L_{2}$, SSIM, LPIPS) measures.
(a) Our model. {$L_{2}$, SSIM, LPIPS} === {.14, .65, .34} (PhotoDNA),
=== {.15, .60, .36} (PDQ), {.20, .40, .50} (NeuralHash)
(b) Our model. {$L_{2}$, SSIM, LPIPS} === {.17, .63, .34} (PhotoDNA),
=== {.21, .50, .34} (PDQ), {.20, .47, .45} (NeuralHash)
(c) Compare images generated by our model and [12].
Figure 5: Samples of true images and reconstructed images generated by hash inversion attacks from hashes of PhotoDNA, PDQ, and NeuralHash. The numbers are quality metrics {$L_{2}$, SSIM, LPIPS} of reconstructed images.
V Conclusions
This paper presents a thorough examination of the security of three real-world perceptual hashing algorithms (PHAs): PhotoDNA, PDQ, and NeuralHash, against three distinct types of attacks: standard image editing attacks, malicious adversarial blackbox attacks, and hash inversion attacks. To support the security study, a query-efficient blackbox adversarial attack algorithm named Joint Soft-label and Hard-label Attack (JSHA) is designed to expedite blackbox adversarial attacks while requiring fewer queries. Additionally, the paper proposes a new data-efficient hash inversion network capable of reliably reconstructing original images from all hashes, even those of very short length.
The findings of the study indicate that PHAs exhibit robustness against blackbox adversarial attacks when realistic constraints regarding adversarial image distortion and query budget are taken into consideration. However, the vulnerabilities to certain image editing methods and the potential information leakage through the hashes raise significant security concerns. These insights underscore the importance of further research and development efforts to enhance the resilience and security of perceptual hashing algorithms in practical applications.

### v2

@@
-Assessing the Adversarial Security of Practical Perceptual Hashing Algorithms
+Robustness of Practical Perceptual Hashing Algorithms to Hash-Evasion and Hash-Inversion Attacks
 Abstract
-Perceptual hashing algorithms (PHAs) are utilized extensively for identifying illegal online content.
-Given their crucial role in sensitive applications, understanding their security strengths and weaknesses is critical.
-This paper compares three major PHAs deployed widely in practice: PhotoDNA, PDQ, and NeuralHash, and assesses their robustness against three typical attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks.
-Contrary to prevailing studies, this paper reveals that these PHAs exhibit resilience to black-box adversarial attacks when realistic constraints regarding the distortion and query budget are applied, attributed to the unique property of random hash variations.
-Moreover, this paper illustrates that original images can be reconstructed from the hash bits, raising significant privacy concerns.
-By comprehensively exposing their security vulnerabilities, this paper contributes to the ongoing efforts aimed at enhancing the security of PHAs for effective deployment.
-Index Terms:
-Perceptual hashing, security, privacy, blackbox adversarial attack, adversarial machine learning
-I Introduction
+Perceptual hashing algorithms (PHAs) are widely used for identifying illegal online content and are thus integral to various sensitive applications.
+However, due to their hasty deployment in real-world scenarios, their adversarial security has not been thoroughly evaluated.
+This paper assesses the security of three widely utilized PHAs—PhotoDNA, PDQ, and NeuralHash—against hash-evasion and hash-inversion attacks.
+Contrary to existing literature, our findings indicate that these PHAs demonstrate significant robustness against such attacks.
+We provide an explanation for these differing results, highlighting that the inherent robustness is partially due to the random hash variations characteristic of PHAs.
+Additionally, we propose a defense method that enhances security by intentionally introducing perturbations into the hashes.
+1 Introduction
 Perceptual hashing, also known as robust hashing, is a content fingerprinting technique that generates similar binary hashes for perceptually similar contents and different hashes for different contents [1].
@@
 Numerous perceptual hash algorithms (PHAs) have been developed or deployed for practical applications [2].
-An early example is the algorithm developed by Google and used on YouTube to identify copyrighted material [3].
-Other examples include those used in online image search engines of Google, Microsoft Bing, etc [4].
-PHAs have gained critical importance, particularly in responding to the reporting of illicit content.
+Early examples include the algorithm used on YouTube to identify copyrighted material [3] and algorithms used in online image search engines of Google, Microsoft Bing, etc [4].
+Recently, PHAs have gained more critical importance, particularly in responding to the requiremnt of reporting illicit content.
 For instance, in cases involving Child Sexual Abuse Media (CSAM), US media service providers are legally obliged to examine suspected illegal content and report it to a “Cyber Tip Line” operated by the National Center for Missing and Exploited Children (NCMEC).
-Microsoft’s PhotoDNA and Meta’s PDQ have been developed for this purpose [5][6][7].
+Microsoft’s PhotoDNA and Meta’s PDQ were immediately deployed for this purpose [5][6][7].
 With a growing emphasis on user privacy, technologies like end-to-end encryption have become prevalent.
@@
 For example, Apple has proposed to deploy NeuralHash on all its iOS devices [9].
-However, deploying PHAs on the client side to scan client data has raised significant concerns regarding client data privacy, prompting widespread criticism about the robustness and privacy-preserving capabilities of the PHAs [10].
-Especially, recent studies have almost unanimously criticized PHAs for their vulnerability to adversarial attacks and the potential leakage of information through hash bits [4][11][12].
-Given their critical role in sensitive applications, it is imperative to thoroughly understand the security of PHAs.
-However, even for PHAs that have been widely deployed in the real world, their security remains insufficiently evaluated.
-Particularly concerning are new PHAs that are hastily deployed without adequate scrutiny.
-To address this gap, this paper systematically evaluates the adversarial robustness and privacy-preserving capabilities of three widely used PHAs: PhotoDNA, PDQ, and NeuralHash.
-To support this study, two new efficient attack algorithms are developed.
+However, this client-side deployment has raised significant concerns among the client users about their data privacy, which inadvertently arouses a widespread criticism regarding the security and privacy-preserving capabilities of PHAs [10].
+Criticism has been further fueled by recent studies highlighting the vulnerability of PHAs to adversarial attacks and the potential for information leakage through hash bits [4][11][12].
+Given the critical role PHAs play in sensitive applications, it is essential to ensure their security.
+Unfortunately, PHAs are often deployed hastily without sufficient security evaluation.
+This paper systematically assesses the adversarial robustness of three widely used PHAs—PhotoDNA, PDQ, and NeuralHash—against two types of attacks: hash-evasion attacks, aimed at enabling illicit content to evade detection, and hash-inversion attacks, aimed at reconstructing the original secret content from hashes.
+To support this evaluation, we introduce two new attack algorithms: a query-efficient black-box adversarial attack algorithm and a data-efficient hash inversion algorithm.
 The major contributions of this paper are as follows.
-  • A systematic comparison of the security and privacy-preserving capability of three major PHAs against three types of attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks, with the development of a query-efficient adversarial attack algorithm and a data-efficient hash inversion network.
-  • Demonstrating the robustness of PHAs against malicious adversarial attacks when realistic constraints over distortion and query budget are imposed, contrary to previous studies that neglected such constraints and thus criticized PHAs as being not robust.
-The robustness is shown to be attributed to the unique property of random hash variations.
-  • Revealing that the hash bits of all three PHAs leak significant information about the content of the original images, allowing for reliable reconstruction of the original image if sufficient context information is available.
-The paper is organized as follows: Section II provides a literature review, Section III formulates the attacks and explains the robustness, Section IV presents the experiment results, and Section V concludes the paper.
-The source code of the experiments, including the binary executable of the PHAs, can be accessed at https://github.com/neddamj/nnhash.
-II Literature Review
-II-A Perceptual Hashing Algorithms
-Among the vast amount of PHAs [2], this paper considers the following three major ones that have been deployed and used widely in practice:
-PhotoDNA, released by Microsoft in 2009, has been used by over 70 companies including Cloudflare and Dropbox [6][13][14].
-To create a hash, PhotoDNA breaks down an image into small pieces, analyzes each piece to extract its unique feature, and converts the features into hash bits.
-PhotoDNA generates a $144$-byte or $1152$-bit hash value for each input image.
-PDQ, which stands for Perceptual algorithm utilising a Discrete Cosine Transform and outputting (amongst others) a Quality metric, was released by Meta in 2019 [7].
-Inspired by pHash [15], PDQ uses the discrete cosine transform (DCT) to extract image features and outputs both a 256-bit hash value and a quality factor.
-NeuralHash was released by Apple in 2021 with the intent to identify CSAM [9].
-An image is first passed through a deep neural network (DNN) to produce a feature embedding.
-The embedding is then converted to a $96$-bit hash.
-The DNN is trained contrastively to ensure that the cosine similarity between perceptually similar images is high while the similarity of perceptually different images is low.
-II-B Adversarial Machine Learning
-The security of PHAs is commonly studied using the terminologies and algorithms of adversarial machine learning [16].
-Specifically, Adversarial robustness measures the resilience of the machine learning models to adversarial attacks.
-Adversarial attacks can be classified as untargeted attacks and targeted attacks.
-In an untargeted attack, the objective is to slightly alter the input to prompt the model to produce a completely different output.
-Conversely, a targeted attack aims to slightly modify the input to elicit a specifically designed output.
-According to the knowledge that the adversary has about the machine learning model, attacks can be categorized to whitebox attacks, where the adversary possesses complete knowledge of the model [16][17][18], and blackbox attacks, where the adversary only has access to the input and output of the model [19][20][21][22].
-In black-box attacks, if the model’s outputs are continuous values like logits or probabilities, the attack is referred to as soft-label attack [19][20][21].
-Otherwise, if the model’s outputs are solely hard-decision labels, the attack is known as a hard-label attack [22].
-II-C Adversarial Robustness of PHAs
-The adversarial robustness of conventional PHAs like pHash [2] was investigated in [23] under the whitebox setting, whereas the robustness under the blackbox setting was explored in [4].
-Many of these studies suggested that PHAs were not robust to adversarial attacks.
-The adversarial robustness of PhotoDNA, PDQ, and NeuralHash has not been adequately addressed, partly because they are new and partly because their source code is not publicly available.
-Regarding their robustness to normal image editing, [24] examined PhotoDNA, while [11] investigated PDQ and NeuralHash.
-Both studies demonstrated that PHAs were robust to certain edits like compression but vulnerable to others like cropping.
-As far as we know, [25] was the first to study these PHAs’ robustness to adversarial attacks.
-It studied NeuralHash under the whitebox setting only, and showed that NeuralHash was not robust to whitebox attacks.
-[26] worked on PDQ and used the linearity of DCT to get adversarial images in a white-box setting.
-[27] generated adversarial images through linear interpolation between two given images.
-Regarding blackbox attacks, [26] studied PDQ and claimed that it was not robust to untargeted blackbox attacks.
-[28] investigated PhotoDNA and PDQ, concluding that they were not robust to either targeted or untargeted attacks.
-These two studies closely resemble the current paper, yet this work differs from them in two key aspects.
-Firstly, previous works computed attack success rates without incorporating limitations on the adversary’s tolerance to image quality degradation or query budget.
-Such constraints are crucial in real-world scenarios.
-It is trivial that unconstrained blackbox attacks will always be successful, particularly in the untargeted setting.
-Secondly, none of the prior works conducted a comparative study of the three PHAs, nor did they investigate the challenge of hash inversion.
-II-D Information Leakage from Hashes
-Regarding privacy preservation, upon its release, Microsoft asserted that the hashes generated by PhotoDNA were irreversible and incapable of reconstructing the original images [29].
-However, in a blog post [12], it was demonstrated that PhotoDNA hashes leak sufficient information about the original image to allow images to be synthesized from the hashes.
-To our knowledge, [12] is the sole study on reconstructing the original image from the hash, albeit focusing solely on PhotoDNA that has long hashes.
-It is unclear if PHAs with short hash suffer from this problem.
-Another approach to investigate information leakage is to train a classifier to classify the image objects from the hash.
-[25][27] illustrated that NeuralHash hashes indeed revealed information about their source images by enabling the identification of the object’s class.
-Interestingly, these findings contradict those of [30] which examined PhotoDNA hashes and claimed that PhotoDNA was resistant to such classification attacks.
-III Attack Algorithms
-Consider a perceptual hash system where the service provider scans the client’s unencrypted content, extracts a hash for each image using a PHA, compares the hash with a database of known illicit hashes, and raises an alarm or forwards the unencrypted content for further analysis in case of hash match [8].
-In this scenario, we consider two adversaries: the client adversary seeking to evade hash matching and the service provider (or some third party) adversary seeking to reconstruct the client’s original images from collected hashes.
-We evaluate the security of PhotoDNA, PDQ, and NeuralHash against three types of attacks: benign image editing attacks and malicious adversarial blackbox attacks employed by the client adversary, and hash inversion attacks employed by the service provider adversary.
-Considering the client adversary, we assess the security of PHAs by measuring the adversary’s attack success rate, where a higher success rate indicates weaker security or robustness.
-Considering the service provider adversary, security is assessed through image reconstruction quality, where higher quality implies weaker security or privacy.
-III-A Image Editing Attacks
-Assume the original input to the PHA is an image ${\bf I}_{0}$, and the true output is an $N$-bit hash ${\bf h}_{0}={\cal H}({\bf I}_{0})$, where ${\cal H}(\cdot)$ denotes the hashing algorithm.
-The client adversary applies standard image editing operations over ${\bf I}_{0}$ to obtain an adversarial image ${\bf I}_{\rm adv}$ with a new $N$-bit hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$.
-A successful attack means ${\bf h}_{\rm adv}$ is sufficiently different from ${\bf h}_{0}$.
-While prior works such as [11] and [24] have investigated image editing attacks on PHAs, a new evaluation is still necessary due to ongoing algorithm updates.
-Notably, NeuralHash has received at least one update since its first release.
-This study addresses the gap in knowledge by evaluating the security of the updated algorithms and, more importantly, by providing the first comparison of the three PHAs.
-III-B Adversarial Blackbox attacks
-III-B1 A Query-efficient Blackbox Attack Algorithm
-This work focuses solely on blackbox, not whitebox, adversarial attacks, reflecting the practical scenario where PHA implementations are proprietary.
-For some PHAs, details regarding architecture and model weights are not publicly available due to their commercial nature.
-For instance, Microsoft enforces non-disclosure agreements for PhotoDNA users, and only an emulation is accessible.
-Similarly, while Ygvar et al.
-(2021) released weights for the first-generation NeuralHash, Apple has since updated the algorithm, keeping the new model confidential.
-Existing blackbox attack algorithms, designed to operate on logits/probabilities or singular class decisions, cannot be directly applied to PHAs because PHAs output a limited number of hash bits, which are neither logits nor class labels.
-Instead, we propose a joint soft-label hard-label attack (JSHA) algorithm, where a soft-label attack step is applied first and its output is passed as input to a hard-label attack step.
-The objective of the soft-label attack step is to modify the image to change the hash without considering the image quality constraint, whereas the objective of the hard-label attack step is to reduce adversarial image quality degradation.
-Consider the original image ${\bf I}_{0}$ with hash value ${\bf h}_{0}={\cal H}({\bf I}_{0})$.
-The objective of the client adversary is to generate an adversarial image ${\bf I}_{\rm adv}$ with new hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$.
-In an untargeted attack, ${\bf h}_{0}$ and ${\bf h}_{\rm adv}$ should be different enough.
-In a targeted attack, there is a target hash ${\bf h}_{\rm tar}\neq{\bf h}_{0}$, which may be obtained from some target image as ${\bf h}_{\rm tar}={\cal H}({\bf I}_{\rm tar})$, and we require that ${\bf h}_{\rm tar}$ and ${\bf h}_{\rm adv}$ are similar.
-In both attacks, the distance between ${\bf I}_{0}$ and ${\bf I}_{\rm adv}$ should be small.
-To simplify notation, let us use normalized Hamming distance to evaluate the similarity of two $N$-bit hashes as
-$D_{h}({\bf h}_{0},{\bf h}_{\rm adv})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}({\bf h}_{0}(n)={\bf h}_{\rm adv}(n))$ (1)
-for two hashes with hash bits ${\bf h}_{0}(n)$ and ${\bf h}_{\rm adv}(n)$, where $\mathbb{I}(x)$ is the indicator function, i.e.
-$\mathbb{I}(x)=1$ if $x$ is true and 00 otherwise.
-In addition, we use per-pixel normalized $L_{2}$ norm to evaluate the distance of two images, i.e.
-$L_{2}({\bf I}_{0},{\bf I}_{\rm adv})=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left({\bf I}_{0}(m)-{\bf I}_{\rm adv}(m)\right)^{2}}$ (2)
-for two images with pixels ${\bf I}_{0}(m)$ and ${\bf I}_{\rm adv}(m)$. Other measures such as SSIM are also used in our experiments.
-The first step of the proposed JSHA algorithm is to apply soft-label blackbox attack to generate an initial adversarial image ${\bf I}_{\rm init}$, with the objective to make the hash ${\bf h}_{\rm init}={\cal H}({\bf I}_{\rm init})$ far away from ${\bf h}_{0}$ for untargeted attacks or close to ${\bf h}_{\rm tar}$ for targeted attacks.
+  • This study demonstrates that PHAs exhibit notable robustness against blackbox adversarial evasion attacks when realistic constraints on image distortion, query budget, and hash difference are applied.
+This finding challenges previous studies that overlooked these constraints, leading to unjustified criticisms of PHAs.
+  • This paper confirms the sufficient privacy-preserving capabilities of hash bits, countering claims from some existing studies that suggested hash bits disclose significant information about the content of the original images.
+The key point is how diverse the images are.
+  • This research reveals that the security of PHAs is partially due to the unique property of random hash variations.
+Based on this observation, a defense method is proposed to enhance security by intentionally randomizing hash bits.
+The paper is organized as follows: Section 2 provides a literature review, Section 3 formulates the attack and defense algorithms, Section 4 presents the experiment results, and Section 5 concludes the paper.
+More detailed experiment data can be found in the Appendix.
+The source code of the experiments, including ways to access binary executables of the PHAs, can be found at https://github.com/neddamj/nnhash.
+2 Literature Review
+Among the numerous PHA’s [2], this paper focuses on the following three major ones due to their widespread deployment in practice:
+  • PhotoDNA, released by Microsoft in 2009, is used by over 70 companies including Cloudflare and Dropbox [6][13][14]. It generates an 1152-bit hash for each input image;
+  • PDQ, launched by Meta in 2019 [7] and heavily inspired by pHash [2][15], produces a 256-bit hash for each input image.
+  • NeuralHash, released by Apple in 2021 with the goal of identifying CSAM [9], creates a more concise 96-bit hash for each input image.
+The adversarial robustness of PhotoDNA, PDQ, and NeuralHash to either hash-evasion or hash-inversion attacks has not been thoroughly investigated, primarily due to their relative novelty and the absence of publicly available source code.
+For hash-evasion, [16] was the first to explore the robustness of NeuralHash, revealing its vulnerability to whitebox attacks, assuming the adversary has full access to the PHA.
+Additionally, [17] examined PDQ and demonstrated how adversarial images could be generated in the whitebox context.
+Furthermore, [18] utilized linear interpolation between two images to create adversarial examples.
+Regarding blackbox hash-evasion attacks, where the adversary only has access to the input and output of the algorithm, [17] examined PDQ.
+It concluded that it was not robust against untargeted blackbox attacks.
+Similarly, [19] evaluated both PhotoDNA and PDQ, finding that they lacked robustness against both targeted and untargeted attacks.
+While these studies are closely related to the current paper, they overlooked the importance of adversaries’ constraints on image distortion, query budget, as well as hash difference levels.
+By all means, unconstrained blackbox attacks may always be successful, particularly in the untargeted setting.
+Regarding hash-inversion attacks or privacy-preserving capabilities of PHAs, Microsoft initially claimed that the hashes generated by PhotoDNA were irreversible and could not be used to reconstruct the original images [20].
+However, a blog post [12] challenged this assertion, suggesting that PhotoDNA hashes might contain enough information to synthesize images from the hashes.
+To the best of our knowledge, [12] is the only work addressing the reconstruction of original images from hashes, and it specifically focuses on PhotoDNA, which uses long hashes.
+Another method for investigating information leakage involves training classifiers to identify image content from hashes.
+Studies such as [16] and [18] demonstrated that NeuralHash hashes could indeed reveal information about their source images, such as identifying the object’s class.
+These findings contrast with those of [21], which examined PhotoDNA hashes and concluded that they were resistant to such classification attacks.
+3 Attack Algorithms and Defense Method
+In a perceptual hash system, a service provider scans a client’s unencrypted content, extracts hashes for each image using a perceptual hashing algorithm (PHA) and compares these hashes with a database of known illicit hashes.
+If a match is found, an alarm is raised or the unencrypted content is forwarded for further analysis [8].
+In this context, we consider two types of adversaries: the client adversary, who aims to evade hash matching, and the service provider (or third-party) adversary, who seeks to reconstruct the client’s original images from collected hashes.
+This paper evaluates the security of three widely used PHAs—PhotoDNA, PDQ, and NeuralHash—against two specific types of attacks: blackbox hash-evasion attacks, which are employed by the client adversary, and hash-inversion attacks, which are utilized by the service provider adversary.
+Additionally, we propose a defense mechanism that can be integrated into the perceptual hashing system.
+3.1 A Query-Efficient Adversarial Blackbox Attack Algorithm for Hash Evasion
+This study focuses exclusively on blackbox adversarial attacks, rather than whitebox attacks, to reflect the practical scenario where PHA implementations are proprietary.
+In many cases, the details of PHA architectures and model weights are not publicly available due to their commercial nature.
+For example, Microsoft enforces non-disclosure agreements for PhotoDNA users, providing only an emulation of the algorithm.
+Similarly, although [22] released weights for the first-generation NeuralHash, Apple has since updated the algorithm, and the details of the latest version remain confidential.
+Consider the original image ${\bf I}_{0}$ with the hash value ${\bf h}_{0}={\cal H}({\bf I}_{0})$.
+The objective of the client adversary is to generate a very similar image ${\bf I}_{\rm adv}$ whose new hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$ is sufficiently different from ${\bf h}_{0}$ in untargeted attacks or sufficiently similar to some target hash ${\bf h}_{\rm tar}$ in targeted attacks.
+We use normalized Hamming distance $D_{h}({\bf h}_{0},{\bf h}_{\rm adv})$ to evaluate the similarity of two hashes, and use per-pixel normalized $L_{2}$ norm $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ to evaluate the similarity of two images.
+Because the output of PHAs is hash bits, not logits or class decisions, we propose a joint soft-label hard-label attack (JSHA) algorithm that can drastically improve blackbox attack query efficiency.
+JSHA consists of two steps.
+The first step is to apply a soft-label blackbox attack to generate an initial adversarial image ${\bf I}_{\rm init}$ so that its hash ${\bf h}_{\rm init}={\cal H}({\bf I}_{\rm init})$ is far away from ${\bf h}_{0}$ for untargeted attacks or close to ${\bf h}_{\rm tar}$ for targeted attacks.
 This step is formulated as the optimization
-$\max_{\delta}\;\;D_{h}\left({\bf h}_{0},{\cal H}({\bf I}_{0}+\delta)\right)$ (3)
-in case of untargeted attacks, or
-$\min_{\delta}\;\;D_{h}\left({\bf h}_{\rm tar},{\cal H}({\bf I}_{0}+\delta)\right)$ (4)
-in case of targeted attacks.
+[…]
 The optimization is conducted iteratively.
-During each iteration, the adversary queries the PHAs with its adversarial input ${\bf I}_{0}+\delta$ to get the hash, and uses the hash to optimize the loss (3) or (4).
-After reaching the query number threshold or the hash distance threshold, this first step stops with output ${\bf I}_{\rm init}={\bf I}_{0}+\delta$.
+In each iteration, the adversary queries the PHAs with its adversarial input ${\bf I}_{0}+\delta$ to get the hash, and uses the hash to optimize the loss.
+After reaching the query number threshold or the hash distance threshold $D_{0}$, this first step stops with output ${\bf I}_{\rm init}={\bf I}_{0}+\delta$.
 The second step applies a hard-label blackbox attack to reduce the image distortion while maintaining the hash distance obtained in the first step.
 Specifically, this step solves the following optimization
-$\min\limits_{{\bf I}_{\rm adv}}\;\;L_{2}({\bf I}_{0},{\bf I}_{\rm adv}),\;\;\;\;\hskip 128.0374pt$ (5)
 […]
-[…]
-where $p$ is the pre-set hash distance threshold.
-Starting from ${\bf I}_{\rm adv}={\bf I}_{\rm init}$, the adversary queries the PHAs to get the hash and uses the queried hash to optimize (5) until the pre-set maximum query number or the threshold of $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ is reached.
-The proposed algorithm is outlined in Algorithm 1.
-One of its major advantages is query efficiency.
-A large learning rate can be applied in the first step to get an initial adversarial sample quickly using a small number of queries.
-In the second step, since we start from ${\bf I}_{\rm init}$ which is only a noisy version of ${\bf I}_{0}$, the convergence is also fast.
-This makes our algorithm able to generate adversarial samples quickly with a very small number of queries.
-Algorithm 1 JSHA: Joint Soft-Label and Hard-Label Attack
-1:Initialization: Original input ${\bf I}_{0}$ and its hash ${\bf h}_{0}$; Target hash ${\bf h}_{\rm tar}$ in case of targeted attacks.
-2:Step One: Unconstrained soft-label attack
-  • The adversary uses queries to optimize (3) or (4)
-  • Stop and output ${\bf I}_{\rm init}$ when the pre-set hash distance threshold or query budget is reached.
-3:Step Two: Constrained hard-label attack
-  • The adversary uses queries to optimize (5)
-  • Stop and output ${\bf I}_{\rm adv}$ when the pre-set image distortion threshold or query budget is reached.
+Starting from ${\bf I}_{\rm adv}={\bf I}_{\rm init}$, the adversary queries the PHAs to get the hash and uses the queried hash to optimize (2) until the query number threshold or the threshold of $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ is reached.
+One of major advantages of this algorithm is high query efficiency.
+The first step is an unconstrained optimization and a large learning rate can be applied to get an initial adversarial sample quickly.
+In the second step, since we start from ${\bf I}_{\rm init}$ which is only a noisy version of ${\bf I}_{0}$, the convergence is fast.
 Refer to caption                   Refer to caption
 (a)                                (b)
-Figure 1: (a) Hash inversion attack model. (b) Architecture of residual block.
-III-B2 Robustness to Adversarial Blackbox Attacks
-Algorithm 1 and most existing blackbox attack algorithms rely on gradient estimation to solve the optimization problem.
-Drawing on the property of PHAs and our prior work on adversarial robust analysis [31], we have the following proposition.
-Proposition 1. Gradient estimation based on queried hash is not reliable.
-Proof.
+Figure 1: (a) Hash inversion attack model (left: for Celeb Dataset. Right: for STL-10 dataset). (b) Architecture of residual block.
+3.2 A Data-Efficient Hash-Inversion Attack Algorithm
+GAN (generative adversarial network) was trained in [12] to invert CelebA images.
+However, the adversary may need more data-efficient hash-inversion attack algorithms because the training data may not be abundant.
+For this, we has developed new and more data-efficient generative models to reconstruct the original input images from the hashes.
+We designed a decoder-style deep model, similar to the Fast Style Transfer [23], to convert the $N$-bit hashes to $(64,64)$ images.
+This model is mainly a combination of residual blocks depicted in Fig.
+1(b), with fractionally-strided convolutions/deconvolutions for learned upsampling.
+The entire architecture is described in Fig.
+1(a), where the left network is for the MNIST and CelebA dataset while the right network is used for the STL-10 dataset.
+For a given training dataset, we first compute the hashes of each image offline and store the image-hash pairs.
+We then train the model outlined above to reconstruct the images as closely as possible by minimizing the mean-square-error (MSE) loss between the model output and the image from which the hash was generated.
+3.3 A Defense Method based on Random Hash Perturbation
+The adversaries rely on gradient estimation to solve the optimization problems.
+It is well known that hash bits of PHAs change slightly and randomly with respect to slight changes in input images.
+Based on existing work on random perturbation defense [24][25], we have the following proposition:
+Proposition 1. Gradient estimation based on queried hash is not reliable due to the random variation of hash bits.
+An outline of the proof is in the Appendix A.1.
+The random variation of hash bits makes the adversaries’ optimization struggle to converge, which will be demonstrated in the next section.
+Furthermore, considering this observation, we propose a defense method where the PHAs randomly invert $q$-portion of the hash bits to guarantee sufficient hash randomness, where $0\leq q\leq 1$.
+4 Experiment
+4.1 Performance Metrics
+We use the attack success rate for assessing the robustness against blackbox adversarial hash-evasion attacks.
+Given a set of $M$ original images ${\bf I}_{0}^{m}$ with their corresponding hashes ${\bf h}_{0}^{m}$, the adversary generates adversarial images ${\bf I}_{\rm adv}^{m}$ with hashes ${\bf h}_{\rm adv}^{m}$.
+The attack success rate is defined as
+${\rm ASR}(p)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{0}^{m},{\bf h}_{\rm adv}^{m})\gtrless p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (3)
+[…]
+For the thresholds, we set $\theta=0.1$.
+On the other hand, we should apply large enough $p$ to evaluate untargeted attacks because ${\bf h}_{\rm adv}$ should be different from ${\bf h}_{0}$.
+It was found in [11] that most pairs of distinct images had $D_{h}({\bf h}_{0},{\bf h}_{1})>0.3$.
+Therefore, we primarily use $p=0.3$ in (3) to determine the success of untargeted attacks.
+In contrast, we should use small enough $p$ for targeted attacks to make sure ${\bf h}_{\rm adv}$ is sufficiently close to ${\bf h}_{\rm tar}$.
+Prior studies [11] indicate that many normal image edits would change an image’s hashes to the level of $D_{h}({\bf h}_{0},{\bf h}_{1})<0.1$.
+Therefore, we mainly use $p=0.1$ in (3) for assessing the success of targeted attacks.
+This important parameter $p$ was unfortunately neglected by most existing adversarial robustness studies.
+For hash-inversion attacks, we use the $L_{2}$ distortion, the SSIM (Structural Similarity Index Measure) [26], and the LPIPS (Learned Perceptual Image Patch Similarity) [27] between the reconstructed image and the true image to evaluate the performance.
+No ASR or query budget is involved.
+4.2 Experiment on Hash-Evasion Attack and Defense
+In implementing our proposed JSHA algorithm, we modified three different soft-label attack algorithms: SimBA [28], NES [29], and ZOsignSGD [30], to use in Step One, and modified the hard-label attack algorithm HSJA [31] to use in Step Two.
+For instance, “SimBA+HSJA” means our JSHA with a modified SimBA as Step One and a modified HSJA as Step Two.
+Modifications included removing/changing constraints and adjusting hyper-parameters, etc.
+We compared the proposed JSHA with the direct application of the original SimBA, NES, and ZOsignSGD algorithms.
+We also compared it with existing works [17] and [19].
+The experiments were conducted on 100 randomly selected images from the ImageNet validation dataset.
+Step One was executed for a total of 3000 queries unless the early stopping criterion was met.
+Step Two was run for an additional 2000 queries unless meeting the early stopping criterion.
+The main challenge for us was the CPU-based non-parallel PHA executables, which limited the number of images or queries we could process.
+In comparison, [19] only utilized 30 images in experiments.
+For targeted hash-evasion attacks, experiment results are presented in Table 1 and Appendix A.3.
+With a focus on ASR(0.1), we observed that all targeted attacks failed, yielding ASR$(0.1)=0$.
+To our knowledge, no literature has ever reported non-zero ASR$(0.1)$, suggesting that all the PHAs are robust against blackbox targeted hash-evasion attacks.
+Table 1: ASR($p$) of Targeted Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+ZO- ZOsign
+SimBA NES+ sign- -SGD
+ASR SimBA +HSJA NES HSJA SGD +HSJA
+PH ASR(.1) 0% 0% 0% 0% 0% 0%
+PD ASR(.1) 0% 0% 0% 0% 0% 0%
+NH ASR(.1) 0% 0% 0% 0% 0% 0%
+[19] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
+[19] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
+For untargeted attacks, experiment results are presented in Table 2 and Appendix A.4.
+With a focus on ASR(0.3), we see that PDQ was robust against all adversarial attacks, with a maximum ASR$(0.3)$ of 1% only.
+It’s worth noting that although [17] achieved a much higher attack success rate against PDQ, it required significantly more queries.
+The results also demonstrated that our defense method was effective because it drastically reduced ASR(0.3) and made the PHAs robust.
+Table 2: Adversary’s ASR($p$) in Untargeted Blackbox Hash-Evasion Attacks, either without defense ($q=0$) or with defense $q=0.1$. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
+ZO- ZOsign Defense Defense
+SimBA NES+ sign- -SGD +SimBA +NES
+ASR SimBA +HSJA NES HSJA SGD +HSJA +HSJA +HSJA
+PH ASR(.3) $1\%$ 92% $0\%$ $69\%$ $0\%$ $19\%$ 27% 21%
+PD ASR(.3) 0% 0% 0% 1% 0% 0% 0% 0%
+NH ASR(.3) 0% 14% 0% 28% 0% 2% 0% 2%
+[19] PH ASR(.3)$=$ 90%, ($L_{2}=0.2$, HSJA with $3\times 10^{4}$ queries)
+[17] PD ASR(.3)$=$ 73%, ($L_{2}=0.1$, up to $8\times 10^{6}$ queries)
+4.3 Experiment on Hash-Inversion Attack and Defense
+Experiments were conducted over three datasets: MNIST, CelebA, and STL-10.
+For each hashing algorithm and dataset, we trained the inversion network depicted in Fig.
+1 for 50 epochs, with a batch size of 64 and an initial learning rate of 0.005.
+We utilized the MSE loss function with the AdamW optimizer and the cosine annealing learning rate scheduler.
+While MNIST and CelebA models were relatively easy to train, we encountered substantial challenges when training STL-10 models.
+We had to degrade STL-10 images to black-and-white in order to get meaningful reconstructed images.
+Some sample images are shown in Fig.
+2 whereas major experiment results are left to Appendix A.5 to save space.
+In general, hash-inversion failed over STL-10 while some level of success could be seen in MNIST and CelebA over PhotoDNA hash only.
+The reason is that the MNIST and CelebA images have similar or regular formation.
+The reconstructed images tend to converge to such common formation when the hash is long enough.
+However, even in this case, the reconstructed CelebA images could not be used to recognize the original face.
+The effectiveness of random hash perturbation defense can be clearly seen in Fig. 2(b). Even the slightest perturbation $q=0.1$ made the reconstructed image quite different from the original image.
+(a) STL-10 image samples with NeuralHash, without defense
+(b) CelebA image samples with PhotoDNA, under various levels $q$ of defense
+Figure 2: Samples of true/original images and reconstructed images by the hash inversion algorithm.
+5 Conclusions
+This paper evaluates the security of three widely used perceptual hashing algorithms (PHAs): PhotoDNA, PDQ, and NeuralHash.
+Contrary to many existing studies, our findings show that these PHAs demonstrate notable robustness against both adversarial blackbox hash-evasion attacks and hash-inversion attacks.
+We attribute this robustness, in part, to the random variations inherent in the hash generation process.
+Building on this insight, we propose enhancing security by introducing additional randomization to the hash bits.
+Appendix A Appendix / supplemental material
+A.1 Proof of Proposition 1 of Section 3.3
 Without loss of generality, consider the optimization problem (1).
@@
 This implies that PHAs are theoretically robust to blackbox attacks under the query budget constraint.
-III-C Hash Inversion Attacks
-GAN (generative adversarial network) was used by [12].
-However, the service provider adversary may need more efficient hash inversion models because the training data may not be abundant in practice.
-For this, we train new and more efficient generative models to reconstruct the original input images from the hashes.
-We designed a decider style DNN, similar to the Fast Style Transfer [32], to convert the $N$-bit hashes to $(64,64)$ images.
-This DNN is mainly a combination of residual blocks, which are described in Fig.
-1(b), with fractionally-strided convolutions/deconvolutions that are used for learned upsampling.
-The entire architecture is described in Fig.
-1(a).
-For a given training dataset, we first compute the hashes of each image offline and store the image-hash pairs.
-We then train the model outlined above to reconstruct the images as closely as possible by minimizing the mean-square-error (MSE) loss between the DNN output and the image from which the hash was generated.
-IV Experiment
-IV-A Performance Metrics
-For assessing the robustness against blackbox attacks, we employ the attack success rate as the evaluation metric.
-Given a set of $M$ original images ${\bf I}_{0}^{m}$ with their corresponding hashes ${\bf h}_{0}^{m}$, the adversary generates adversarial images ${\bf I}_{\rm adv}^{m}$ along with their hashes ${\bf h}_{\rm adv}^{m}$.
-The attack success rate is defined as
-${\rm ASR}(p)=$
-$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{0}^{m},{\bf h}_{\rm adv}^{m})>p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (7)
-for untargeted attacks, and
-${\rm ASR}(p)=$
-$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left(D_{h}({\bf h}_{\rm tar}^{m},{\bf h}_{\rm adv}^{m})<p)\right)\mathbb{I}\left(L_{2}({\bf I}_{0}^{m},{\bf I}_{\rm adv}^{m})<\theta\right)$ (8)
-for targeted attacks (with the target hash ${\bf h}_{\rm tar}$).
-$p\in[0,1]$ is the pre-set hash distance threshold and $\theta$ is the pre-set image distortion threshold.
-The adversary can adjust the threshold $p$ to control the tradeoff between the true positive rate (TPR) and the false positive rate (FPR) of the hash or image similarity decisions.
-Untargeted attacks use a large $p$ because ${\bf h}_{\rm adv}$ should be different from ${\bf h}_{0}$.
-As shown in Fig.
-2, when we calculated $D_{h}({\bf h}_{0},{\bf h}_{1})$ of distinct images of the ImageNet validation dataset, we found that very few pairs of images had $D_{h}({\bf h}_{0},{\bf h}_{1})<0.3$, which was also observed in [11].
-Therefore, we primarily use $p=0.3$ in (7) to determine the success of untargeted attacks.
-In contrast, targeted attacks use a small $p$ to make sure ${\bf h}_{\rm adv}$ is sufficiently close to ${\bf h}_{\rm tar}$.
-Prior studies [11] indicate that many normal image edits would change an image’s hashes to the level of $D_{h}({\bf h}_{0},{\bf h}_{1})<0.1$.
-Therefore, we mainly use $p=0.1$ in (8) for assessing the success of targeted attacks.
-To comprehensively assess attack success, we collect ASR$(p)$ for all $p=0.1$, $0.2$, $0.3$, $0.4$ during experiments. We set the image distortion threshold $\theta=0.1$.
-Figure 2: Cumulative distribution function (cdf) of normalized hamming distance $D_{h}({\bf h}_{0},{\bf h}_{1})$ between hashes of distinct images in the ImageNet validation dataset.
-For hash inversion attacks, we use the $L_{2}$ distortion (2), the SSIM (Structural Similarity Index Measure) [33], and the LPIPS (Learned Perceptual Image Patch Similarity) [34] between the reconstructed image and the true image to evaluate the performance.
-No ASR or query budget is involved.
-IV-B Experiment Results of Image Editing Attacks
+A.2 Robustness Problem of PHAs to Normal Image Editing
+One of the special issue of PHAs is that they seem not robust to some normal imaging editing methods, for which [32] examined PhotoDNA, while [11] investigated PDQ and NeuralHash.
+Both studies demonstrated that PHAs were robust to certain edits like compression but vulnerable to others like cropping.
+While such prior works have already extensively investigated image editing on PHAs, a new evaluation is necessary due to ongoing algorithm updates.
+Notably, NeuralHash has received at least one update since its first release.
+This study addresses the gap in knowledge by evaluating the robustness of the updated algorithms and, more importantly, by providing the first comparison of the three PHAs.
+Assume the original input to the PHA is an image ${\bf I}_{0}$, and the true output is an $N$-bit hash ${\bf h}_{0}={\cal H}({\bf I}_{0})$, where ${\cal H}(\cdot)$ denotes the hashing algorithm.
+The client adversary applies standard image editing operations over ${\bf I}_{0}$ to obtain an adversarial image ${\bf I}_{\rm adv}$ with a new $N$-bit hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$.
+A successful attack means ${\bf h}_{\rm adv}$ is sufficiently different from ${\bf h}_{0}$.
 We experimented with four image editing operations that can keep the original high image quality: JPEG compression with random quality factors, resizing images with random scaling factors, filtering images with the vignette filter, or rotating images to random degrees.
@@
 Using 1000 images from the ImageNet validation dataset, we obtained ASR$(p)$ for various $p$ based on (3), wherein the $L_{2}$ condition was removed by setting $\theta$ to infinity.
-Experiment results are presented in Tables I.
+Experiment results are presented in Tables 3.
 The data for ASR(.3) is highlighted in black because $p=0.3$ is the threshold we suggested to focus on.
@@
 NeuralHash was the most robust among the three.
-Surprisingly, the newest edition of NeuralHash that we experimented with was still not robust to filtering, rotating, and mirroring, similar to the first edition experimented in [11][25], even though DNNs can be trained to be invariant to such standard operations with data augmentation.
-TABLE I: Adversary’s ASR($p$) with Image Editing Attacks
+Surprisingly, the newest edition of NeuralHash that we experimented with was still not robust to filtering, rotating, and mirroring, similar to the first edition experimented in [11][16].
+Nevertheless, we believe this robustness problem can be mitigated by training the model to be invariant to such standard operations with data augmentations.
+Table 3: Adversary’s ASR($p$) with Image Editing Attacks
 ASR Compress Resize Filter Rotate
@@
 NeuralHash ASR(.4) 0% 0% 12% 31%
-IV-C Experiment Results of Untargeted Blackbox Attacks
-In implementing our proposed JSHA algorithm, we modified three different soft-label attack algorithms: SimBA [21], NES [19], and ZOsignSGD [20], to use in Step One, and modified the hard-label attack algorithm HSJA [22] to use in Step Two.
-For instance, “SimBA+HSJA” means our JSHA with a modified SimBA as Step One and a modified HSJA as Step Two.
-Modifications included removing/changing constraints and adjusting hyper-parameters, etc.
-Specifically, in Step One, we skipped image distortion constraints and adopted extremely large perturbation step sizes, e.g, using a perturbation magnitude of $\epsilon=0.9$ for SimBA instead of the original $\epsilon=0.01$ as described in [21].
-The experiments were conducted on 100 randomly selected images from the ImageNet validation dataset.
-Step One was executed for a total of 3000 queries unless the early stopping criterion was met.
-Step Two was run for an additional 2000 queries unless meeting the early stopping criterion.
-The main challenge for us was the CPU-based non-parallel PHA executables, which limited the number of images or queries we could process.
-In comparison, [28] only utilized 30 images in experiments.
-Experiment results are presented in Tables II.
-We compared the proposed JSHA with the straightforward application of standard SimBA [21], NES [19], and ZOsignSGG [20] algorithms.
-We also compared it with [26] and [28].
-With a focus on ASR(0.3), the data indicated that NeuralHash was relatively robust because the maximum ASR$(0.3)$ was $28\%$ only.
-PDQ was robust against all adversarial attacks, with a maximum ASR$(0.3)$ of 1% only.
-It’s worth noting that although [26] achieved a much higher attack success rate against PDQ, it required significantly more queries.
-It could afford millions of queries due to its access to the PDQ source code, enabling faster computation.
-Furthermore, the experiment data revealed that our proposed JSHA outperformed standard blackbox attack algorithms.
-SimBA+HSJA surpassed [28] in attacking PhotoDNA, as the latter produced more noisy adversarial images and utilized more queries.
+A.3 Extra Experiment Data of Targeted Hash-Evasion Attacks in Section 4.2
+Besides the data shown in Section 4.2 and Table 1, more experiment data of targeted attacks are presented in Table 4.
+We find that the robustness is extremely strong because of ASR$(p)\approx 0$ across all $p$.
+The only exception was NeuralHash, which exhibited ASR(0.4) of 43% with SimBA.
+However, this isn’t indicative of successful adversarial attacks; rather, it’s due to numerous hashes generated by NeuralHash having $D_{h}({\bf h}_{0},{\bf h}_{1})$ as low as 0.3.
+Under NeuralHash, many images are inherently adversarial when evaluated at $p=0.4$.
+Therefore, contrary to the prior claims that PHAs were not robust against targeted attacks [19], our findings demonstrate the strong robustness of all PHAs against targeted blackbox attacks.
+This is significant to the practical application of PHAs because the client adversary needs targeted attacks to ensure that its adversarial hashes do not inadvertently match illicit hashes in the database.
+The strong robustness means it is hard for the adversary to undertake either untargeted or targeted attacks without being caught.
+Table 4: Adversary’s ASR($p$) in Targeted Blackbox Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+ZO- ZOsign
+SimBA NES+ sign- -SGD
+ASR SimBA +HSJA NES HSJA SGD +HSJA
+PH ASR(.1) 0% 0% 0% 0% 0% 0%
+PH ASR(.2) 0% 0% 0% 0% 0% 0%
+PH ASR(.3) 0% 0% 0% 0% 0% 0%
+PH ASR(.4) 1% 2% 0% 0% 0% 0%
+PD ASR(.1) 0% 0% 0% 0% 0% 0%
+PD ASR(.2) 0% 0% 0% 0% 0% 0%
+PD ASR(.3) 0% 0% 0% 0% 0% 0%
+PD ASR(.4) 0% 0% 0% 0% 0% 0%
+NH ASR(.1) 0% 0% 0% 0% 0% 0%
+NH ASR(.2) 0% 0% 0% 0% 0% 0%
+NH ASR(.3) 0% 0% 0% 0% 0% 0%
+NH ASR(.4) 43% 11% 0% 0% 0% 0%
+[19] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
+[19] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
+A.4 Extra Experiment Data of Untargeted Blackbox Hash-Evasion Attacks in Section 4.2
+A.4.1 Experiment data and sample images of untargeted attacks without defense
+Besides the experiment data shown in Section 4.2 and Table 2, extra experiment data are presented in Table 5.
 Sample adversarial images generated with our proposed JSHA with various pre-set threshold $p$ are depicted in Fig.
-3, with their corresponding distortion metrics listed in Table III.
-Notably, the three images with $p=0.4$ were deemed invalid adversarial images as their distortion was not under the threshold $\theta=0.1$, resulting in extremely noisy images.
+3 and Fig.
+4, with their corresponding distortion metrics listed in Table 6.
+Notably, the images with $p=0.4$ were deemed invalid adversarial images as their distortion was not under the threshold $\theta=0.1$, resulting in extremely noisy images.
 Similarly, for PDQ, the adversarial images with $p=0.2$ and $0.3$ were also invalid.
-In summary, contrary to previous studies suggesting that PHAs were not robust against untargeted attacks, we found that some PHAs like PDQ and NeuralHash could be considered as robust if realistic constraints regarding image distortion and query budget are taken into account.
-TABLE II: Adversary’s ASR($p$) with Untargeted Black-box Attacks.
-(PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
+Therefore, contrary to previous studies suggesting that PHAs were not robust against untargeted attacks, we found that some PHAs like PDQ and NeuralHash could be considered as robust if realistic constraints regarding image distortion and query budget are taken into account.
+Table 5: Adversary’s ASR($p$) in Untargeted Blackbox Hash-Evasion Attacks. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
 ZO- ZOsign
@@
 [17] PD ASR(.3)$=$ 73%, ($L_{2}=0.1$, up to $8\times 10^{6}$ queries)
-(a) Original image ${\bf I}_{0}$.
-PhotoDNA                  PDQ                   NeuralHash
-(b) Adversarial images ${\bf I}_{\rm adv}$.
-Figure 3: Samples of an original image and the adversarial images created by the proposed untargeted blackbox attack algorithm JSHA (NES+HSJA).
-TABLE III: Distortion of the Adversarial Image Samples in Fig. 3(b).
+Table 6: Distortion of the Adversarial Image Samples in Fig. 3(b).
 Measure Threshold $p$ PhotoDNA PDQ NeuralHash
@@
 $p=0.4$ 0.17 0.19 0.19
-IV-D Experiment Results of Targeted Blackbox Attacks
-The experiment setting of targeted attacks was similar to that of untargeted attacks, except for utilizing ${\bf h}_{\rm tar}$ obtained from random images instead of ${\bf h}_{0}$.
-Results are presented in Tables IV.
-With a focus on ASR(0.1) in this case, we observed that all targeted attacks failed, yielding ASR$(0.1)=0$.
-To our knowledge, no literature has ever reported non-zero ASR$(0.1)$, suggesting that all the PHAs are robust against targeted attacks.
-The robustness is extremely strong because of ASR$(p)\approx 0$ across all $p$.
-The only exception was NeuralHash, which exhibited ASR(0.4) of 43% with SimBA.
-However, this isn’t indicative of successful adversarial attacks; rather, it’s due to numerous hashes generated by NeuralHash having $D_{h}({\bf h}_{0},{\bf h}_{1})$ as low as 0.3, as illustrated in Fig.
-2.
-Under NeuralHash, many images are inherently adversarial when evaluated at $p=0.4$.
-In summary, contrary to the prior claims that PHAs were not robust against targeted attacks [28], our findings demonstrate the strong robustness of all PHAs against targeted blackbox attacks.
-This is significant to the practical application of PHAs because the client adversary needs targeted attacks to ensure that its adversarial hashes do not inadvertently match illicit hashes in the database.
-The strong robustness means it is hard for the adversary to undertake either untargeted or targeted attacks without being caught.
-TABLE IV: Adversary’s ASR($p$) with Targeted Blackbox Attacks.
-(PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
-ZO- ZOsign
-SimBA NES+ sign- -SGD
-ASR SimBA +HSJA NES HSJA SGD +HSJA
-PH ASR(.1) 0% 0% 0% 0% 0% 0%
-PH ASR(.2) 0% 0% 0% 0% 0% 0%
-PH ASR(.3) 0% 0% 0% 0% 0% 0%
-PH ASR(.4) 1% 2% 0% 0% 0% 0%
-PD ASR(.1) 0% 0% 0% 0% 0% 0%
-PD ASR(.2) 0% 0% 0% 0% 0% 0%
-PD ASR(.3) 0% 0% 0% 0% 0% 0%
-PD ASR(.4) 0% 0% 0% 0% 0% 0%
-NH ASR(.1) 0% 0% 0% 0% 0% 0%
-NH ASR(.2) 0% 0% 0% 0% 0% 0%
-NH ASR(.3) 0% 0% 0% 0% 0% 0%
-NH ASR(.4) 43% 11% 0% 0% 0% 0%
-[28] PH ASR(.3) ($L_{2}=0.19$) 83% ($2\times 10^{6}$ queries)
-[28] PD ASR(.3) ($L_{2}=0.42$) 100% ($6\times 10^{5}$ queries)
-IV-E Experiment Results of Hash Inversion Attacks
-To demonstrate the service provider adversary’s capability to reconstruct original images from their hashes under some contextual knowledge, such as the group of images to which the original images may belong, we conducted experiments using two datasets MNIST and CelebA as known contexts.
-For each hashing algorithm and dataset, we trained the inversion network depicted in Fig.
-1 for 50 epochs, with a batch size of 64 and an initial learning rate of 0.005.
-We utilized the MSE loss function with the AdamW optimizer and the cosine annealing learning rate scheduler.
-Results are presented in Table V.
-Samples of synthesized images along with their original images can be observed in Fig.
-4 and Fig.
-5.
-It is evident that the hash leaks sufficient information for image reconstruction when certain contextual information is known, such as whether the image belongs to MNIST digits or CelebA photos.
-Although no image quality metrics were given in [12], Fig.
-5(c) shows that our proposed model had a performance comparable to that of [12].
-In Table V, smaller $L_{2}$ and LPIPS mean better image reconstruction, while larger SSIM means better image reconstruction.
-PhotoDNA has the longest hashes at 1152-bits, PDQ has 256-bit hashes and NeuralHash has 96-bit hashes.
-Intuitively, this implies that PhotoDNA hashes contain the most information about the original images while NeuralHash hashes contain the least.
-Table V supported this intuition.
-The results in the MNIST for PDQ were left blank because PDQ could not work on MNIST.
-PDQ outputs quality scores of 0 for every image, and each of the hash values calculated turned out to be a 256-dimensional vector of zeros for MNIST images.
-TABLE V: Quality of Adversarial Images Created in Hash Inversion Attacks (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+(a) Original image ${\bf I}_{0}$.
+PhotoDNA                        PDQ                        NeuralHash
+(b) Adversarial images ${\bf I}_{\rm adv}$.
+Figure 3: Samples of an original image and the adversarial images created by the proposed untargeted blackbox attack algorithm JSHA (NES+HSJA).
+PhotoDNA                        PDQ                        NeuralHash
+Figure 4: Samples of an original image and the adversarial images created by the proposed untargeted blackbox attack algorithm JSHA (NES+HSJA).
+A.4.2 Experiment data of untargeted attacks under proposed defense
+Observing that some untargeted attack algorithms such as SimBA+HSJA achieved relataively high ASR(0.3) over some PHAs such as PhotoDNA, we experimented with the hash perturbation defense method proposed in Section 3.3.
+The level of our random hash perturbation was denoted as $q$, where $0\leq q\leq 1$, which stand for the ratio of the hash bits that were randomly flipped.
+Besides the experiment results presented in Section 4.2 and Table 2, more experiment data are presented in Tables 7, 8, and 9.
+It can be seen that random perturbation reduced ASR($p$) drastically, which means the attacks became much less successful or even failed.
+Even a slight perturbation with $q=0.1$ could prevent most of the attacks.
+Table 7: Adversary’s ASR($0.1$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
+ZOsign
+SimBA NES+ -SGD
+ASR +HSJA HSJA +HSJA
+PH q =0.1 100% $100\%$ $100\%$
+PH q =0.2 100% $100\%$ $100\%$
+PH q =0.3 100% $100\%$ $100\%$
+PD q =0.1 0% 0% 0%
+PD q =0.2 0% 0% 0%
+PD q =0.3 0% 0% 0%
+NH q =0.1 21% 18% 15%
+NH q =0.2 20% 18% 15%
+NH q =0.3 21% 18% 15%
+Table 8: Adversary’s ASR($0.2$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
+ZOsign
+SimBA NES+ -SGD
+ASR +HSJA HSJA +HSJA
+PH q =0.1 98% 99% 95%
+PH q =0.2 98% 99% 95%
+PH q =0.3 98% 99% 95%
+PD q =0.1 0% 0% 0%
+PD q =0.2 0% 0% 0%
+PD q =0.3 0% 0% 0%
+NH q =0.1 5% 5% 2%
+NH q =0.2 5% 6% 3%
+NH q =0.3 5% 6% 2%
+Table 9: Adversary’s ASR($0.3$) in Untargeted Blackbox Hash-Evasion Attacks when Defense was Applied. (PH: PhotoDNA. PD: PDQ. NH: NeuralHash. )
+ZOsign
+SimBA NES+ -SGD
+ASR +HSJA HSJA +HSJA
+PH q =0.1 27% 21% 5%
+PH q =0.2 27% 21% 5%
+PH q =0.3 27% 21% 5%
+PD q =0.1 0 0% 0%
+PD q =0.2 0% 0% 0%
+PD q =0.3 0% 0% 0%
+NH q =0.1 0% 2% 0%
+NH q =0.2 0% 2% 0%
+NH q =0.3 0% 0% 0%
+A.5 Extra Experiment Results of Hash-Inversion Attacks in Section 4.3
+A.5.1 Experiment data and sample images of hash-inversion attacks without defense
+Besides the experiment sample images presented in Section 4.3 and Fig.
+2, a complete set of experiment data of image reconstruction quality are in Table 10.
+In addition, Fig.
+5 shows sample images of MNIST dataset reconstruction, Fig.
+6 shows sample images of STL-10 dataset reconstruction, whereas Fig.
+7 shows sample images of CelebA dataset reconstruction.
+All these experiment results were obtained without the defense being applied.
+Generally, the hash-inversion attacks worked with some success when there was a lot of regularity in the dataset (e.g., MNIST and CelebA).
+In this case, the hash-inversion algorithm gave an image of such regularity only.
+The longer hash values leaked more information and allowed better hash-inversion.
+Nevertheless, even in the best case, it was still impossible to match the reconstructed human faces with the true faces in the CelebA images.
+When the dataset was diverse (e.g., STL-10), we were not able to learn any meaningful inversions.
+This demonstrated that the PHAs were robust against hash-inversion attacks in practical applications.
+In addition, PHAs with smaller number of output bits were always preferable because natural images were not very standardized, so the smaller number of bits could help to prevent information leakage/hash inversion.
+The reason for [12] to draw the false conclusion of successful hash-inversion was because only CelebA images were used over the long PhotoDNA hashes, and the authors overlooked the fact that the faces can not be discriminated based on the reconstructed images.
+Table 10: Quality of Adversarial Images Created in Hash Inversion Attacks (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
 Dataset Hash $L_{2}$ SSIM LPIPS
@@
 NH $0.23\pm 0.05$ $0.29\pm 0.09$ $0.53\pm 0.07$
+PH $\mathbf{0.14}\pm 0.03$ $\mathbf{0.41}\pm 0.09$ $\mathbf{0.56}\pm 0.06$
+STL-10 PD $0.24\pm 0.05$ $0.26\pm 0.05$ $0.65\pm 0.02$
+NH $0.23\pm 0.05$ $0.14\pm 0.09$ $0.67\pm 0.07$
 (a) PhotoDNA. Left: (.14, .83, .05). Right: (.12, .86, .07)
 (b) NeuralHash. Left: (.26, .61, .10). Right: (.34, .52, .14)
-Figure 4: Samples of true images and adversarial images generated by hash inversion attacks based on hashes of (a) PhotoDNA, and (b) NeuralHash. The numbers are ($L_{2}$, SSIM, LPIPS) measures.
+Figure 5: Samples of true MNIST images and adversarial images generated by hash inversion attacks based on hashes of (a) PhotoDNA, and (b) NeuralHash. The numbers are ($L_{2}$, SSIM, LPIPS) measures.
+(a) Using PhotoDNA Hash bits
+(b) Using PDQ Hash bits
+(c) Using NeuralHash hash bits
+Figure 6: Samples of true STL-10 images and adversarial images generated by hash inversion attacks.
 (a) Our model. {$L_{2}$, SSIM, LPIPS} === {.14, .65, .34} (PhotoDNA),
@@
 (c) Compare images generated by our model and [12].
-Figure 5: Samples of true images and reconstructed images generated by hash inversion attacks from hashes of PhotoDNA, PDQ, and NeuralHash.
+Figure 7: Samples of true CelebA images and reconstructed images generated by hash inversion attacks from hashes of PhotoDNA, PDQ, and NeuralHash.
 The numbers are quality metrics {$L_{2}$, SSIM, LPIPS} of reconstructed images.
-V Conclusions
-This paper presents a thorough examination of the security of three real-world perceptual hashing algorithms (PHAs): PhotoDNA, PDQ, and NeuralHash, against three distinct types of attacks: standard image editing attacks, malicious adversarial blackbox attacks, and hash inversion attacks.
-To support the security study, a query-efficient blackbox adversarial attack algorithm named Joint Soft-label and Hard-label Attack (JSHA) is designed to expedite blackbox adversarial attacks while requiring fewer queries.
-Additionally, the paper proposes a new data-efficient hash inversion network capable of reliably reconstructing original images from all hashes, even those of very short length.
-The findings of the study indicate that PHAs exhibit robustness against blackbox adversarial attacks when realistic constraints regarding adversarial image distortion and query budget are taken into consideration.
-However, the vulnerabilities to certain image editing methods and the potential information leakage through the hashes raise significant security concerns.
-These insights underscore the importance of further research and development efforts to enhance the resilience and security of perceptual hashing algorithms in practical applications.
+A.5.2 Experiment data hash-inversion attacks under the proposed defense
+To mitigate the potential hash inversion attacks over regular images, we experimented with the proposed random hash perturbation defense algorithm.
+Some sample images were shown in Fig.
+2(b) and explained briefly in Section 4.3.
+Extra experiment data are shown in Tables 11, 12 and 13.
+Compare these data with those in Table 10, it is clear that as we increased the level of perturbation to the hash bits, the quality of reconstructed images decreased as expected.
+Therefore, the random perturbation defense does help to defend against hash-inversion attacks.
+Table 11: Quality of Inverted Images when Defended with q = 0.1 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+Dataset Hash $L_{2}$ SSIM LPIPS
+PH $0.14\pm 0.03$ $0.41\pm 0.00$ $0.56\pm 0.06$
+STL-10 PD $0.24\pm 0.05$ $0.26\pm 0.05$ $0.65\pm 0.02$
+NH $0.23\pm 0.05$ $0.14\pm 0.07$ $0.67\pm 0.03$
+Table 12: Quality of Inverted Images when Defended with q = 0.2 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+Dataset Hash $L_{2}$ SSIM LPIPS
+PH $0.18\pm 0.04$ $0.26\pm 0.08$ $0.68\pm 0.05$
+STL-10 PD $0.23\pm 0.05$ $0.24\pm 0.08$ $0.65\pm 0.04$
+NH $0.23\pm 0.05$ $0.17\pm 0.08$ $0.77\pm 0.07$
+Table 13: Quality of Inverted Images when Defended with q = 0.3 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
+Dataset Hash $L_{2}$ SSIM LPIPS
+PH $0.20\pm 0.06$ $0.21\pm 0.09$ $0.71\pm 0.05$
+STL-10 PD $0.23\pm 0.06$ $0.20\pm 0.08$ $0.68\pm 0.04$
+NH $0.23\pm 0.05$ $0.17\pm 0.08$ $0.78\pm 0.04$